OpenStudy (anonymous):

A computer downloads files at a constant rate. The table shows how many megabytes the computer downloads over specific lengths of time. What is the slope of the line that represents the computer downloads? @myininaya Computer Downloads Minutes 3 5 7 9 Megabytes 11.25 18.75 26.25 33.75

OpenStudy (anonymous):

where is the table?

OpenStudy (anonymous):

Oh sorry wait

OpenStudy (anonymous):

Computer Downloads Minutes 3 5 7 9 Megabytes 11.25 18.75 26.25 33.75 @M0j0jojo

OpenStudy (anonymous):

@TheSmartOne can you help me with this question and then a few more :S

TheSmartOne (thesmartone):

11.25/3=3.75 18.75/5=3.75 26.25/7=3.75 33.75/9=3.75

TheSmartOne (thesmartone):

So if y is megabytes and x is minutes the equation would be y=3.75x

TheSmartOne (thesmartone):

So your slope is 3.75

OpenStudy (anonymous):

Ah! Can you help with 4 more c: I think I get it but I really need to pass the quiz

OpenStudy (anonymous):

@TheSmartOne Can you help me with 4 more??? :c